OpenStudy (anonymous):

Is 4, 15, 26, 37, 48, 59 - arithmetic or geometric sequence or neither?

OpenStudy (anonymous):

arithemetic

OpenStudy (anonymous):

Arithmetic because 15-4= 26-15= 37-26 = 48-37= 11 which is common difference.

OpenStudy (anonymous):

Thank You for explaining. : D